# Efficient Feature Extraction and Classification Architecture for MRI-Based Brain Tumor Detection

### v1

@@ Abstract
Uncontrolled cell division in the brain is what gives rise to brain tumors. If the tumor size increases by more than half, there is little hope for the patient’s recovery. This emphasizes the need of rapid and precise brain tumor diagnosis. When it comes to analyzing, diagnosing, and planning therapy for brain tumors, MRI imaging plays a crucial role. A brain tumor’s development history is crucial information for doctors to have. When it comes to distinguishing between human soft tissues, MRI scans are superior. In order to get reliable classification results from MRI scans quickly, deep learning is one of the most practical methods. Early human illness diagnosis has been demonstrated to be more accurate when deep learning methods are used. In the case of diagnosing a brain tumor, when even a little misdiagnosis might have serious consequences, accuracy is especially important. Disclosure of brain tumors in medical images is still a difficult task. Brain MRIs are notoriously imprecise in revealing the presence or absence of tumors. Using MRI scans of the brain, a Convolutional Neural Network (CNN) was trained to identify the presence of a tumor in this research. Results from the CNN model showed an accuracy of 99.17%. The CNN model’s characteristics were also retrieved. In order to evaluate the CNN model’s capability for processing images, we applied the features via the following machine learning models: KNN, Logistic regression, SVM, Random Forest, Naive Bayes, and Perception. CNN and machine learning models were also evaluated using the standard metrics of Precision, Recall, Specificity, and F1 score. The significance of the doctor’s diagnosis enhanced the accuracy of the CNN model’s assistance in identifying the existence of tumor and treating the patient.
@@ I Introduction
Brain is a core part of CNS (Central Nervous System), which regulates all physiological and cognitive activities i.e. thought, emotion, touch, motor skils, vision, respiration etc. [1]. Brain tumor is uncontrolled cell division in brain or other parts of the CNS that causes malfunctioning. The malignancy of a tumor depends on how fast the cell reproduces. Non-malignant (benign and not cancerous) tumors grow slowly and do not spread into other tissues. Malignant brain tumors are cancerous. Most of the time, they multiply and invade neighboring healthy tissues [2] [3]. In USA, there are an estimated 20,500 primary brain tumors identified each year; 3750 of these instances include people under the age of 19 and 2870 involved children under the age of 15 [4]. In comparison to tumors in any other organ of the human body, diagnosing a tumor in the brain is particularly difficult. The blood-brain barrier (BBB) surrounds the brain, making it impossible for regular radioactive markers to detect the tumor cells’ increased activity. Additionally, tumor size, shape, location, and type make early detection more challenging [3] [5]. Brain tumor incident is more common in developed countries. Australia, North America, and Northern Europe have the highest rates of brain tumors. On the other hand, Africa has the lowest rate [6].
Deep learning and machine learning are two new technologies that have significantly developed different fields of applications [7, 8, 9, 10]. In particular, a vast field of study has been opened up in medical image processing, and several studies are currently being conducted in this field of study. An important aspect of this study area involves automating the segmentation and categorization of brain tumors. Akram et. al [11] introduced a computer-aided method for identifying tumors in which they segmented tumors using the global thresholding methodology. Sharjar et al [12] presented an automated method for detecting brain tumor. The photos were preprocessed using image enhancement filters, followed by additional segmentation and the extraction of characteristics to identify the tumor. Parasuraman et al[13] used a feed-forward neural network to classify tumor and normal regions. The method involves four steps: pre-processing with filters, image segmentation with clustering, feature extraction using gray-level co-occurrence matrix (GLCM), and tumor classification with ensemble classifier. Irmak et al. [14] developed three CNN models for three different datasets. Accuracy for the three datasets was 99.33%, 92.66%, and 98.14% respectively. Although multi classification accuracy wasn’t as good as binary classification.
@@ II Materials and Methods
We implemented a CNN model from scratch, three pre-trained models, and five traditional ML models in our proposed method. Our main objective is to diagnose brain tumors effectively and precisely by sending MRI pictures of the tumors to a CNN. Fig. 1 represents the workflow of our study. Labeled MRI images are supplied into a CNN feature extractor after minimum preprocessing, and the retrieved features are fed into a classification layer. Finally, we compared the performance of our trained model to the current state of the art in various aspects.
@@ II-D1 KNN
KNN is a Supervised learning algorithm which can solve both classification and regression predicting problems. It is a lazy algorithm because it uses all of the data for training during classification and lacks a training phase. In this algorithm, the number K of the neighbors has been chosen. Then the Euclidean distance between K neighbors has been calculated and K nearest neighbors were chosen based on distance value. Then, for each category, the number of data points among those K neighbors has been counted. Finally, the category with the highest neighbor count has been given to the new data points.
@@ II-D2 SVM
Parametric classification aims to describe the usual feature space values or distribution of each class. SVM, in contrast, focuses solely on the training samples that are situated closest to the ideal class border in the feature space[19][20]. The name of the approach comes from these samples, which are known as support vectors. The SVM classifier is fundamentally binary in that it recognizes only one distinction between two classes.
@@ II-D3 Naive Bayes
Naive Bayesian networks (NB) are the most basic types of Bayesian networks. They are made up of DAGs with a single parent (representing the invisible node) and a large number of children (corresponding to the visible nodes), and they make a strong assumption that each child node is independent of the other child nodes.As a result, the independence model (Naive Bayes) is based on estimating:
@@ II-D4 Random Forest
Random Forest is a classifier that employs several decision trees on various subsets of the input dataset and averages the outcomes to improve the projected accuracy of the dataset[21]. The final class is determined by majority voting across all of the trees. The trees will be individually less accurate because of the lower training data and variables, but the ensemble as a whole will be more accurate because of the reduced correlation between the trees. Due to the presence of several trees, RF has the specific advantage that no individual trees need to be pruned.
@@ II-D5 Multi Layer Perceptron
This classifier solves a quadratic programming issue with linear constraints as opposed to the non-convex, unconstrained minimization problem that neural networks often use to find their weights[22]. In a multilayer perceptron, neurons are structured in layers and totally coupled with one another by edges to create a directed graph, which is a feedforward artificial neural network. According to the definition of “feedforward”, this graph is acyclic. Each of these edges has an actual value known as the edge weight attached to it. The input layer, a number of hidden layers, and the output layer are the layers of multilayer perceptron neural networks. There is a summation function and an activation function for every neuron. The summation function is expressed as follows:
$\begin{split}S_{j}=\sum_{i=1}^{n}W_{i}{}_{j}I_{i}+\beta_{j}\end{split}$ (2)
The activation function accepts the output of this summation function as an input. The set of weights is adjusted to estimate the desired output after the neural network has been built.
@@ III Results and Discussion
In this paper, we have implemented four different models which are Densenet, Resnet50, EfficientnetB0, and our own CNN model. For optimization of our deep learning model, we used Adam algorithm. First, the performance is evaluated using several performance metrics. During the training of a model, we concentrated on reducing loss while simultaneously boosting accuracy. Table I displays the validation accuracy of each model. Here, we can observe that the validation accuracy was best achieved by our own CNN model. Fig. 7(b) also displays the outcome graphically. Then, using these models, extracted features are sent to various machine learning models. And the outcome is revealed on Fig 7(a). From the Table II, we can conclude that Scratch+SVM, DenseNet+SVM, ResNet+SVM, and EfficientNet+Logistic Regression perform better than any other combination of machine learning and pretrained models and the comparison is illustrated in Fig. 7(a).
@@ III-A Performance metrics
Evaluation metrics are used to measure the quality of a model. Evaluating models or algorithms is essential for any project. A model can be tested using a wide variety of evaluation metrics. Confusion metrics had been applied in this scenario. An N*N matrix, where N is the number of expected classes, is a confusion matrix. Since N=2 applies to our issue, we receive a 2*2 matrix. True Positive (TP) indicates the number of accurately categorized attack records. True Negative (TN) is the proportion of correctly categorized normal records. False Positive (FP) is the quantity of typical records that were misclassified. False negatives (FN) is the number of attack records that were wrongly categorised.
  • Accuracy (AC):
    $\begin{split}AC=\frac{TP+TN}{TP+TN+FP+FN}\end{split}$ (3)
  • Precision (P):
    $\begin{split}P=\frac{TP}{TP+FP}\end{split}$ (4)
  • Recall (R):
    $\begin{split}R=\frac{TP}{TP+FN}\end{split}$ (5)
  • F1 Score (F):
    $\begin{split}F={(\frac{recall^{-1}+precision^{-1}}{2})}^{-1}\end{split}$ (6)
  • Spicificity (SP):
    $\begin{split}SP=\frac{TN}{TN+FP}\end{split}$ (7)
@@
TABLE I: Validation accuracy of CNN models
Our CNN 99.17%
DenseNet 98.5%
ResNet50 99%
EfficientNetB0 98.33%
TABLE II: Accuracy of classifiers for CNN models
@@ III-A1 Comparison with other works
We will now compare our conclusions to a variety of other approaches that were suggested in the literature review. The comparative result is shown in Table III. It covers the findings of some of the newest methods suggested for locating brain tumors. The table demonstrates that WCNN was used to achieve the highest accuracy, which was 99.3. The other methods are less accurate than this. Even if the accuracy of our own CNN model is less accurate than that of the Sarhan et al.[23], it is still more accurate than alternative approaches. A number of pretrained models that we also trained showed impressive accuracy. Then, in order to increase the accuracy of our models, we mix these models with machine learning models. The highest accuracy we can achieve is 99.83%, which is higher than the paper by Sarhan et al.[23].
TABLE III: Performance comparison with other models
@@ IV Future Works and Conclusion
In this paper, we explore different models to detect brain tumor more effectively and precisely. We have surpassed all previously investigated methods with an accuracy of 99.83%. Since brain tumor can lead to cancer, the impact of brain tumors are terrible for us and endanger our lives. We believe that our approach has the potential to reduce the risk of developing cancer and save many lives. In the future, we hope to develop a model that can accurately detect all types of tumors. Although our model solely analyzes MRI dataset, we also wish to take into account other medical imaging methods, such as CT (Computed Tomography) scan, PET (Positron-Emission Tomography) etc.

### v2

@@ Abstract
Uncontrolled cell division in the brain is what gives rise to brain tumors. If the tumor size increases by more than half, there is little hope for the patient’s recovery. This emphasizes the need for rapid and precise brain tumor diagnosis. MRI imaging plays a crucial role in analyzing, diagnosing, and planning therapy for brain tumors. A brain tumor’s development history is crucial information for doctors to have. When it comes to distinguishing between human soft tissues, MRI scans are superior. Deep learning is one of the most practical methods for quickly getting reliable classification results from MRI scans. Early human illness diagnosis has been demonstrated to be more accurate when deep learning methods are used. In the case of diagnosing a brain tumor, when even a little misdiagnosis might have serious consequences, accuracy is especially important. Disclosure of brain tumors in medical images is still a difficult task. Also because of the scarcity of annotated data localizing the tumor regions is quite expensive task. Brain MRIs are notoriously imprecise in revealing the presence or absence of tumors. Using MRI scans of the brain, a Convolutional Neural Network (CNN) was trained to identify the presence of a tumor in this research. Results from the CNN model showed an accuracy of 99.17%. The CNN model’s characteristics were also retrieved and we also localized the tumor regions from the unannotated images using GradCAM, a deep learning explainability tool. In order to evaluate the CNN model’s capability for processing images, we applied the features via the following machine learning models: KNN, Logistic regression, SVM, Random Forest, Naive Bayes, and Perception. CNN and machine learning models were also evaluated using the standard metrics of Precision, Recall, Specificity, and F1 score. The significance of the doctor’s diagnosis enhanced the accuracy of the CNN model’s assistance in identifying the existence of tumor and treating the patient.
@@ I Introduction
Brain is a core part of CNS (Central Nervous System), which regulates all physiological and cognitive activities i.e. thought, emotion, touch, motor skils, vision, respiration etc. [1]. Brain tumor is uncontrolled cell division in brain or other parts of the CNS that causes malfunctioning. The malignancy of a tumor depends on how fast the cell reproduces. Non-malignant (benign and not cancerous) tumors grow slowly and do not spread into other tissues. Malignant brain tumors are cancerous. Most of the time, they multiply and invade neighboring healthy tissues [2] [3]. In USA, there are an estimated 20,500 primary brain tumors identified each year; 3750 of these instances include people under the age of 19 and 2870 involved children under the age of 15 [4]. In comparison to tumors in any other organ of the human body, diagnosing a tumor in the brain is particularly difficult. The blood-brain barrier (BBB) surrounds the brain, making it impossible for regular radioactive markers to detect the tumor cells’ increased activity. Additionally, tumor size, shape, location, and type make early detection more challenging [3, 5].
Deep learning and machine learning are two new technologies that have significantly developed different fields of applications [6, 7, 8, 9]. An important aspect of this study area involves automating the segmentation and categorization of brain tumors. Akram et. al [10] introduced a computer-aided method for identifying tumors in which they segmented tumors using the global thresholding methodology.
Parasuraman et al[11] used a feed-forward neural network to classify tumor and normal regions. The method involves four steps: pre-processing with filters, image segmentation with clustering, feature extraction using gray-level co-occurrence matrix (GLCM), and tumor classification with ensemble classifier. Irmak et al. [12] developed three CNN models for three different datasets. However, training such models and validation requires a significant amount of pixel level annotated data and in the field of medical imaging getting such data is too costly and requires domain experts.
In this study, we proposed a convolution Neural Network based architecture that aims to detect tumorous brain MRI images and localize the affected regions. We achieved an accuracy of 99.83% without pixel level annotated data. Finally we applied the technique proposed in GradCAM [13] for producing visual explanations for decisions from our CNN model. This method localizes the tumorous region from the gradients of the final convolution layer of the model which eliminates the need for pixel level annotation for localizing the anomalous regions.
@@ II Materials and Methods
We implemented a CNN model from scratch, three pre-trained models, and five traditional ML models in our proposed method. Our main objective is to diagnose brain tumors effectively and precisely by sending MRI pictures of the tumors to a CNN and localize the anomalous regions. Fig. 1 represents the workflow of our study. Labeled MRI images are supplied into a CNN feature extractor after minimum preprocessing, and the retrieved features are fed into a classification layer. Finally, we compared the performance of our trained model to the current state of the art in various aspects.
@@ II-D1 KNN
KNN is a Supervised learning algorithm which can solve both classification and regression-predicting problems. It is a lazy algorithm because it uses all of the data for training during classification and lacks a training phase.
@@ II-D2 SVM
SVM focuses solely on the training samples that are situated closest to the ideal class border in the feature space. The name of the approach comes from these samples, which are known as support vectors. The SVM classifier is fundamentally binary in that it recognizes only one distinction between two classes.
@@ II-D3 Naive Bayes
Naive Bayesian networks (NB) are made up of DAGs with a single parent (representing the invisible node) and a large number of children (corresponding to the visible nodes), and they make a strong assumption that each child node is independent of the other child nodes. As a result, the independence model is based on estimating:
@@ II-D4 Random Forest
Random Forest is a classifier that employs several decision trees on various subsets of the input dataset and averages the outcomes to improve the projected accuracy of the dataset. The final class is determined by majority voting across all of the trees.
@@ II-D5 Multi Layer Perceptron
This classifier solves a quadratic programming issue with linear constraints as opposed to the non-convex, unconstrained minimization problem that neural networks often use to find their weights. In a multilayer perceptron, neurons are structured in layers and totally coupled with one another by edges to create a directed graph. The input layer, a number of hidden layers, and the output layer are the layers of multilayer perceptron neural networks. There is a summation function and an activation function for every neuron.
@@ III Results and Discussion
In this research study, we have implemented four different models which are Densenet, Resnet50, EfficientnetB0, and our own CNN model. For optimization of our deep learning model, we used Adam algorithm. First, the performance is evaluated using several performance metrics. During the training of a model, we concentrated on reducing loss while simultaneously boosting accuracy. Table LABEL:tab:Accuracy-CNN displays the validation accuracy of each model. Here, we can observe that the validation accuracy was best achieved by our own CNN model. Fig. 8(b) also displays the outcome graphically. Then, using these models, extracted features are sent to various machine learning models. And the outcome is revealed on Fig 8(a). From the Table I, we can conclude that Scratch+SVM, DenseNet+SVM, ResNet+SVM, and EfficientNet+Logistic Regression perform better than any other combination of machine learning and pretrained models and the comparison is illustrated in Fig. 8(a).
Figure 5: Localization of tumorous region in Brain MRI
@@ III-A Performance metrics
Evaluation metrics are used to measure the quality of a model. A model can be tested using a wide variety of evaluation metrics. Confusion metrics is an N*N matrix, where N is the number of expected classes, is a confusion matrix. Since N=2 applies to our issue, we receive a 2*2 matrix. True Positive (TP) indicates the number of accurately categorized attack records. True Negative (TN) is the proportion of correctly categorized normal records. False Positive (FP) is the quantity of typical records that were misclassified. False negatives (FN) are the number of attack records that were wrongly categorized.
[…]
@@
TABLE I: Accuracy of classifiers for CNN models
@@ III-A1 Comparison with other works
We will now compare our conclusions to a variety of other approaches that were suggested in the literature review. The comparative result is shown in Table II. It covers the findings of some of the newest methods suggested for locating brain tumors. The table demonstrates that WCNN was used to achieve the highest accuracy, which was 99.3. The other methods are less accurate than this. Even if the accuracy of our own CNN model is less accurate than that of the Sarhan et al.[18], it is still more accurate than alternative approaches. A number of pretrained models that we also trained showed impressive accuracy. Then, in order to increase the accuracy of our models, we mix these models with machine learning models. The highest accuracy we can achieve is 99.83%, which is higher than the paper by Sarhan et al.[18].
TABLE II: Performance comparison with other models
@@
III-B Anomalous region localization
We utilized the GradCAM method for localizing the tumorous regions of the images. The Gradient-weighted Class Activation Mapping (Grad-CAM) approach utilizes the gradients of a specified target class, which propagate into the final convolutional layer, to generate a coarse localization map that emphasizes significant areas in the image relevant to the class’s prediction. Figure 5 shows the model’s robust output for the localization technique. The sample images are not annotated pixel by pixel. Only the labels are used to generate the localization heatmap for the images.
@@ IV Future Works and Conclusion
In this study, we explored different models to detect brain tumor more effectively and precisely. We have surpassed all previously investigated methods with an accuracy of 99.83%. We also proposed a tumor localization method which localizes tumorous regions without pixel level annotation. The robust localization of the regions and the accurate detection illuminates the path to use lightweight models for anomaly detection in medical images. Since brain tumor can lead to cancer, the impact of brain tumors are terrible for us and endanger our lives. We believe that our approach has the potential to reduce the risk of developing cancer and save many lives. In the future, we hope to develop a model that can accurately detect all types of tumors. Although our model solely analyzes MRI dataset, we also wish to take into account other medical imaging methods, such as CT (Computed Tomography) scan, PET (Positron-Emission Tomography) etc.